A medication is 75% effective against a bacterial infection. Find the probability that if 12 people take the medication, at least 1 person's infection will not improve
Answer by stanbon(75887)

A medication is 75% effective against a bacterial infection. Find the probability that if 12 people take the medication, at least 1 person's infection will not improve
-------
Binomial Problem with n = 12 and p(effective) = 0.75 ; P(not) = 0.25
----
P(1<= x <=12 not improve) = 1 - P(x = 0 improve) = 1 - 0.25^12 = 0.999999940400..
------------------------
Cheers,
Stan H.